# GBR: Generative Bundle Refinement for High-fidelity Gaussian Splatting and Meshing

### v1

@@ VII Conclusion
In this paper, we presented GBR: Generative Bundle Refinement, a framework designed for high-fidelity Gaussian splatting and meshing. By integrating neural bundle adjustment, generative depth refinement, and a multimodal loss function, GBR significantly improves geometric accuracy, detail preservation, and robustness in Gaussian splatting optimization. Our method effectively recovers camera parameters, generates dense point maps, and supports novel view synthesis and mesh reconstruction from sparse input images. We validated the performance of GBR on the DTU, TNT, and MIP datasets, demonstrating its superiority over state-of-the-art methods in both geometric accuracy and novel view rendering quality under sparse-view conditions. Finally, we demonstrated GBR’s applicability to large-scale real-world scenes, such as the Pavilion of Prince Teng and the Great Wall.

### v2

@@ VII Conclusion
In this paper, we presented GBR: Generative Bundle Refinement, a framework designed for high-fidelity Gaussian splatting and meshing. By integrating neural bundle adjustment, generative depth refinement, and a multimodal loss function, GBR significantly improves geometric accuracy, detail preservation, and robustness in Gaussian splatting optimization. Our method effectively recovers camera parameters, generates dense point maps, and supports novel view synthesis and mesh reconstruction from sparse input images. We validated the performance of GBR on the DTU, Tanks & Temples, and Mip-NeRF 360 datasets, demonstrating its superiority over state-of-the-art methods in both geometric accuracy and novel view rendering quality under sparse-view conditions. Finally, we demonstrated GBR’s applicability to large-scale real-world scenes, such as the Pavilion of Prince Teng and the Great Wall.